# Best Arm Identification for Prompt Learning under a Limited Budget

### v2

@@ Abstract
The remarkable instruction-following capability of large language models (LLMs) has sparked a growing interest in automatically learning suitable prompts. However, while many effective methods have been proposed, the cost incurred during the learning process (e.g., accessing LLM and evaluating the responses) has not been considered. To overcome this limitation, this work explicitly incorporates a finite budget constraint into prompt learning. Towards developing principled solutions, a novel connection is established between prompt learning and fixed-budget best arm identification (BAI-FB) in multi-armed bandits (MAB). Based on this connection, a general framework TRIPLE (besT aRm Identification for Prompt LEarning) is proposed to harness the power of BAI-FB in prompt learning systematically. Unique characteristics of prompt learning further lead to two embedding-based enhancements of TRIPLE by exploiting the ideas of clustering and function approximation. Extensive experiments on multiple well-adopted tasks using both GPT 3.5 and Llama2 demonstrate the significant performance improvement of TRIPLE over the previous baselines while satisfying the limited budget constraints. Code available at https://github.com/ShenGroup/TRIPLE
@@ 1 Introduction
Large language models (LLMs) have rapidly changed technology landscapes in our society (Devlin et al.,, 2018; Touvron et al., 2023a, ; Touvron et al., 2023b, ; Bubeck et al.,, 2023; OpenAI, 2023a, ). Researchers continuously find effective ways to unlock their potential on various downstream tasks. Among different research directions, the remarkable ability of LLMs to follow instructions has motivated the study of searching for suitable prompts to interact with them (Liu et al.,, 2023). This approach is particularly attractive as it does not require updating the inside parameters of an LLM, and is natural in the way of human conversations. One representative example is the success of zero-shot Chain-of-Thoughts prompting (Kojima et al.,, 2022), i.e., adding a seemingly simple phrase of “let’s think step by step” bumps the reasoning capability of LLMs. Nevertheless, it has also been recognized that the performance of an LLM is sensitive to the selected prompts (Zhao et al.,, 2021; Lu et al.,, 2021), and manually designing suitable prompts can be a labor-intensive process (Mishra et al.,, 2021). Thus, there is a growing interest to automatically perform prompt learning, e.g., Zhou et al., (2022); Xu et al., (2022); Diao et al., (2022); Deng et al., (2022); Prasad et al., (2022); Zhang et al., 2023a ; Guo et al., (2023); Pan et al., (2023); Pryzant et al., (2023); Yang et al., (2023).
@@
While these studies on prompt learning have gradually moved towards more practical scenarios of finding discrete and interpretable prompts with only black-box LLM accesses, the costs incurred during learning have not been considered. In particular, most of the previous works do not measure the required number of accesses to LLMs. However, in practice, accesses to LLMs are often (1) financially costly (e.g., each OpenAI API access incurs a cost); (2) time-wise consuming (e.g., even a locally hosted LLM would typically require seconds to respond); (3) under total usage limits (e.g., OpenAI has hard per-day and per-month limits on API accesses). Furthermore, it is often overlooked that evaluating the responses of an LLM for different prompts can be costly as many prompting tasks (e.g., writing improvement, mathematical reasoning, etc) would require human (and sometimes domain expert) opinions.
@@
To address these challenges, this work proposes to study prompt learning under an explicitly imposed budget constraint when interacting with the targeted LLM, in addition to the previously considered constraints (e.g., discrete, interpretable, and black-box). To the best of our knowledge, budget constraints are only briefly mentioned in Zhou et al., (2022); Pryzant et al., (2023), and there are no systematic or principled investigations of how to address the limited budget constraint in prompt learning. The main contributions of this work are summarized as follows.
@@
$\bullet$ The constraint of a limited budget is explicitly introduced into prompt learning, which has been largely ignored before. As most of the existing methods rely on selecting from a pre-generated candidate prompt pool, we focus our study on how to carefully allocate budgets to test each candidate prompt so that the optimal one can be learned efficiently and effectively.
@@
$\bullet$ We develop a general solution framework, termed TRIPLE (besT aRm Identification for Prompt LEarning), by establishing a novel connection between prompt learning and multi-armed bandits (MAB) (Lattimore and Szepesvári,, 2020). In particular, we focus on harnessing the power of fixed-budget best arm identification (BAI-FB) (Audibert et al.,, 2010; Karnin et al.,, 2013) to address prompt learning (especially, selection) with a limited budget constraint. Two representative designs TRIPLE-SH and TRIPLE-CR, inspired by celebrated BAI-FB algorithms, are presented.
$\bullet$ To better handle large prompt pools and improve the scalability of our design, two enhanced methods, TRIPLE-CLST and TRIPLE-GSE, are proposed, where prompt embeddings are leveraged to accelerate learning. TRIPLE-CLST clusters prompts based on their embeddings for a two-phased scheme to first find the optimal cluster. TRILE-GSE, on the other hand, incorporates the idea of function approximation, i.e., sharing information among prompts via a learned function over their embeddings.
@@
$\bullet$ Extensive experimental results are reported using well-adopted prompt tasks to demonstrate the superiority of the TRIPLE solutions over previous baselines. In particular, when handling small prompt pools, the basic TRIPLE-SH and TRIPLE-CR already outperform all prior baselines by (on average) $13\%$ to $35\%$; when tackling large prompt pools, the enhanced TRIPLE-CLST and TRIPLE-GSE further achieve additional $13\%$ to $16\%$ average gain over TRIPLE-CR, all on GPT 3.5. The gains are even more pronounced on Llama2-7b. Moreover, the proposed methods can be directly plugged into two popular prompt learning pipelines, APE (Zhou et al.,, 2022) and APO (Pryzant et al.,, 2023), with end-to-end performances significantly improved over their original implementations.
Besides the aforementioned technical contributions, we believe the value of this work extends broadly to providing a new perspective of prompt learning from MAB, and also a new application scenario of MAB in prompt learning. This established connection may spark further innovations in both fields as discussed in Sec. 7.
2 Prompt Learning under a Limited Budget
Following Zhou et al., (2022); Chen et al., (2023), we present a concrete formulation of the problem of prompt learning.
@@ 2 Prompt Learning under a Limited Budget
Consider that we are using an LLM $f(\cdot)$, which provides a mapping from any input $X\in{\mathcal{V}}$ to a distribution $\Delta_{{\mathcal{V}}}$ over the language space ${\mathcal{V}}$. The answer $\hat{Y}\in{\mathcal{V}}$ given by the LLM is assumed to be sampled from $f(X)$ as $\hat{Y}\sim f(X)$. Note that instead of treating $f(\cdot)$ as a deterministic function providing a specific output answer, we generally consider the practical setting where the answers of LLM exhibit a certain level of randomness.
@@
For prompt learning, we aim to find a prompt $p$ such that when concatenated with inputs $X$ of a certain task (i.e, as $[p;X]$), it provides good performance in expectation with respect to the input distribution ${\mathcal{I}}_{X}$ and the inherent randomness of LLM $f(\cdot)$. The performance is measured as
$\mu(p):=\mathbb{E}_{X\sim{\mathcal{I}}_{X}}\mathbb{E}_{\hat{Y}\sim f([p;X])}\left[s(X,\hat{Y})\right],$
where $s(X,\hat{Y})$ denotes a score function that measures the quality of the output $\hat{Y}$ for the input $X$.
Motivated by the common usage scenario of LLMs, recent studies have imposed several constraints on this learning problem (Zhou et al.,, 2022; Pan et al.,, 2023; Chen et al.,, 2023; Guo et al.,, 2023), where the three key ones are
  • (I)
    Black-box: the method can be applied to black-box LLMs, i.e., only have access to an API $f(\cdot)$ and no access to the intermediate structure or parameters inside (including gradients, output likelihood, etc);
  • (II)
    Discrete: the learned prompt must be discrete characters, instead of continuous values (i.e., soft prompts);
  • (III)
    Interpretable: the learned prompt must be understandable by humans, instead of gibberish words.
@@
Intuitively, the process of learning a good prompt would require interactions with the LLM (i.e., sample $\hat{Y}\sim f([p;X])$ and evaluating its responses (i.e., obtain score $s(X,\hat{Y})$). However, as mentioned in Sec. 1, such interactions and evaluations are often costly. Thus, besides the aforementioned constraints, we further explicitly take into account that the prompt learning process should have a limited budget:
  • (IV)
    Limited budget: the total number of trials with the LLM that happen during the learning is at most $N$.
Finally, the prompt learning problem considered in this work can be formulated as:
Finding $p^{*}$ with high performance $\mu(p^{*})$ under constraints of black-box, discrete, interpretable, and a limited budget.
@@
Directly tackling this prompt learning problem has been widely recognized as challenging even without the constraint of a limited budget (Liu et al.,, 2023). As highlighted in Pryzant et al., (2023); Chen et al., (2023), it essentially requires performing a black-box discrete optimization. Instead, many proposed methods rely on the pipeline of first generating a pool of candidate prompts and then selecting from it (Jiang et al.,, 2020; Zhou et al.,, 2022; Xu et al.,, 2022; Prasad et al.,, 2022; Guo et al.,, 2023). The prompt generation can either be performed manually or follow designed automatic protocols. For example, the famous APE design (Zhou et al.,, 2022) selects from prompts generated by an LLM using demonstrations. From a unified perspective, we can simplify the problem into generating a pool of prompts ${\mathcal{P}}$ and finding the optimal prompt in it:
$p^{*}:=\operatorname*{arg\,max}_{p\in{\mathcal{P}}}\mu(p).$
While many efforts have been devoted along this line, we recognize that these previous methods (which are thoroughly reviewed in Sec. 6) are largely focused on how to generate prompts, while limited attention has been paid to how to select from the already generated prompts. For example, Jiang et al., (2020); Xu et al., (2022); Guo et al., (2023); Prasad et al., (2022) directly evaluate all the generated prompts on the entire development dataset; Zhou et al., (2022) adopts a naive filtering strategy without theoretical or empirical justification. Such treatment is understandable in previous works where budget limitations are not considered, i.e., unlimited evaluations can be performed for selection. With an explicit budget limitation, however, we need to carefully allocate the budgets to each prompt so that the optimal prompt (or at least a sufficiently good one) can be correctly learned, which is the main focus of this work. An overview of the considered prompt learning pipeline and our focus is illustrated in Fig. 1.
@@
Figure 1: An overview of the commonly adopted prompt learning pipeline, which consists of two components: prompt generation and selection. Previous works mostly investigate the prompt generation component and ignore costs incurred during selection, where GrIPS and APE are proposed in Prasad et al., (2022); Zhou et al., (2022). This work, instead, mainly focuses on the prompt selection component, especially with an explicit budget constraint.
3 Connecting Prompt Learning with Best Arm Identification
@@
We provide a new perspective of prompt learning through the lens of tools in multi-armed bandits (MAB) (Lattimore and Szepesvári,, 2020; Lai and Robbins,, 1985). In particular, prompt learning under a limited budget is shown to be intrinsically aligned with the problem of fixed-budget best-arm identification (BAI-FB) (Audibert et al.,, 2010; Karnin et al.,, 2013). In the following, a brief introduction to MAB is first provided. Then the connection between prompt learning (especially selection) with MAB (especially BAI-FB) is established. Based on this connection, we propose to both leverage existing and develop new BAI-FB designs to tackle the problem of prompt learning with a limited budget.
@@ 3.1 Multi-armed Bandits
The research of multi-armed bandits (MAB) has a long and rich history; see representative surveys of Lattimore and Szepesvári, (2020); Bubeck et al., (2012). The most basic form of MAB, i.e., the finite-armed stochastic bandits, considers a system with a set ${\mathcal{K}}$ finite arms (i.e., actions) that provide stochastic rewards when pulled. When interacting with the system, the agent can select one arm $k\in{\mathcal{K}}$ to pull at each time, and she receives a stochastic reward:
$r_{k}\sim\texttt{dist}_{k}(\nu_{k}),$
where $\texttt{dist}_{k}(\nu_{k})$ denotes the action $k$’s reward distribution with an unknown expectation $\nu_{k}$.
The learning objective of the agent in MAB can be roughly divided into two categories: (1) regret minimization, which maximizes the expected cumulative rewards collected by the agent (Auer et al.,, 2002; Audibert et al.,, 2009; Garivier and Cappé,, 2011); (2) best arm identification, which targets at outputting the best arm $k^{*}=\operatorname*{arg\,max}_{k\in{\mathcal{K}}}\nu_{k}$ (Audibert et al.,, 2010; Garivier and Kaufmann,, 2016; Jamieson and Nowak,, 2014). These two objectives often require different learning strategies. Regret minimization typically relies on a carefully designed balance between exploration (i.e., obtaining new information) and exploitation (i.e., collecting higher rewards based on the previous information). Best arm identification, on the other hand, is also called pure exploration as it only focuses on obtaining information to find the best arm.
3.2 A Bandit View of Prompt Learning
Based on the above introduction, it can be intuitively understood that the prompt learning (especially, selection) problem can be mapped into an MAB setting:

### v3

# Efficient Prompt Optimization Through the Lens of Best Arm Identification

@@ Abstract
The remarkable instruction-following capability of large language models (LLMs) has sparked a growing interest in automatically finding good prompts, i.e., prompt optimization. Most existing works follow the scheme of selecting from a pre-generated pool of candidate prompts. However, these designs mainly focus on the generation strategy, while limited attention has been paid to the selection method. Especially, the cost incurred during the selection (e.g., accessing LLM and evaluating the responses) is rarely explicitly considered. To overcome this limitation, this work provides a principled framework, TRIPLE, to efficiently perform prompt selection under an explicit budget constraint. TRIPLE is built on a novel connection established between prompt optimization and fixed-budget best arm identification (BAI-FB) in multi-armed bandits (MAB); thus, it is capable of leveraging the rich toolbox from BAI-FB systematically and also incorporating unique characteristics of prompt optimization. Extensive experiments on multiple well-adopted tasks using various LLMs demonstrate the remarkable performance improvement of TRIPLE over baselines while satisfying the limited budget constraints. As an extension, variants of TRIPLE are proposed to efficiently select examples for few-shot prompts, also achieving superior empirical performance.
@@ 1 Introduction
Large language models (LLMs) have rapidly changed technology landscapes in our society (Devlin et al.,, 2018; Touvron et al., 2023a, ; Touvron et al., 2023b, ; Bubeck et al.,, 2023; OpenAI, 2023a, ). Researchers continuously find effective ways to unlock their potential on various downstream tasks. Among different research directions, the remarkable ability of LLMs to follow instructions has motivated the study of searching for suitable prompts to interact with them (Liu et al.,, 2023). This approach is particularly attractive as it does not require updating the inside parameters of an LLM, and is natural in the way of human conversations. Nevertheless, it has also been recognized that the performance of an LLM is sensitive to the selected prompts (Zhao et al.,, 2021; Lu et al.,, 2021), and manually designing suitable prompts can be a labor-intensive process (Mishra et al.,, 2021). Thus, there is a growing interest to perform automatic prompt optimization (Zhou et al.,, 2022; Xu et al.,, 2022; Diao et al.,, 2022; Deng et al.,, 2022; Prasad et al.,, 2022; Zhang et al., 2023a, ; Guo et al.,, 2023; Pan et al.,, 2023; Pryzant et al.,, 2023; Yang et al.,, 2023).
@@
While these studies have proposed different prompt optimization designs, they commonly follow the approach of generating a pool of candidate prompts and then selecting from them. With a deeper look, it can be recognized that the focus in these existing works largely leans towards how to generate the candidate pool, while limitation attention have been paid towards how to select from the candidates. For example, many works (Jiang et al.,, 2020; Xu et al.,, 2022; Guo et al.,, 2023; Prasad et al.,, 2022) directly evaluate all the generated prompts on the entire development dataset. However, this less-emphasized selection process typically requires accesses to LLMs, which are often (1) financially costly (e.g., each OpenAI API access incurs a cost); (2) time-wise consuming (e.g., even a locally hosted LLM would typically require seconds to respond); (3) under total usage limits (e.g., OpenAI has hard per-day and per-month limits on API accesses). Furthermore, it is often overlooked that evaluating the responses of an LLM for different candidate prompts can be costly as many tasks (e.g., writing improvement, mathematical reasoning, etc.) would require human (and sometimes domain expert) opinions. As a result, the prompt optimization process can incur an unaffordable cost without a proper selection method.
@@
To make the learning process more accessible, this work proposes to study prompt optimization under an explicitly imposed budget constraint when interacting with the targeted LLM, in addition to the previously considered requirements (e.g., discrete, interpretable, and black-box). To the best of our knowledge, budget constraints are only briefly mentioned in Zhou et al., (2022); Pryzant et al., (2023), and there are no systematic or principled investigations of how to address the limited budget constraint in prompt optimization. The main contributions of this work are summarized as follows.
@@
$\bullet$ The constraint of a limited budget is explicitly introduced into prompt optimization, which has been largely ignored before. As most of the prompt optimization methods rely on selecting from a pre-generated candidate prompt pool, we focus our study on how to carefully allocate budgets to test each candidate prompt so that the optimal one can be learned efficiently and effectively.
@@
$\bullet$ We propose a general solution framework, termed TRIPLE (besT aRm Identification for Prompt LEarning), by establishing a novel connection between prompt optimization and multi-armed bandits (MAB) (Lattimore and Szepesvári,, 2020). In particular, we focus on harnessing the power of fixed-budget best arm identification (BAI-FB) (Audibert et al.,, 2010; Karnin et al.,, 2013) to address prompt optimization (especially, selection) with a limited budget constraint. Two representative designs TRIPLE-SH and TRIPLE-CR, inspired by celebrated BAI-FB algorithms, are presented. To improve scalability, two enhanced methods, TRIPLE-CLST and TRIPLE-GSE, are further proposed, where prompt embeddings are leveraged by exploiting the ideas of clustering and function approximation to accelerate the learning process.
@@
$\bullet$ Extensive experimental results are reported using well-adopted prompt tasks and varying LLMs to demonstrate the superiority of TRIPLE over previous baselines. In particular, on GPT3.5 and Llama2, compared with baseline methods also not using prompt embeddings, the basic TRIPLE-SH and TRIPLE-CR achieves performance improvements by (on average) $3\%$ to $16\%$. When leveraging prompt embeddings, the enhanced TRIPLE-CLST and TRIPLE-GSE also outperform corresponding baselines by (on average) $10\%$ to $56\%$ with fewer prompts than budget and (on average) $16\%$ to $45\%$ with more prompts than budget. The gains are further evidenced on other LLMs, i.e., Gemma and Mistral. Moreover, the proposed methods can be directly plugged into two popular prompt optimization pipelines, APE (Zhou et al.,, 2022) and APO (Pryzant et al.,, 2023), with end-to-end performances significantly improved over their original implementations.
$\bullet$ This work extends broadly to providing a new perspective of prompt optimization from MAB, and also a new application scenario of MAB in prompt optimization. This established connection may spark further innovations in both fields. As one concrete example, we extend the study to optimizing the selection of examples in few-shot prompts (Brown et al.,, 2020), which can be recognized as a BAI-FB problem in the setup of combinatorial bandits (Chen et al.,, 2014; Bubeck et al.,, 2013). Experimental results illustrate that the extensions of TRIPLE achieve superior performance, demonstrating its rich potential.
Key Related Works. We discuss a few works that explicitly or implicitly touch upon the selection efficiency in prompt optimization, and a complete literature review can be found in Appendix A. First, Zhou et al., (2022) discusses a naive filtering strategy without theoretical or empirical justifications. Chen et al., (2023) leverages Bayesian optimization (BO) with expected improvement (EI) as the acquisition function to select continuous soft prompts. BO can be viewed as similar to BAI while mostly focusing on infinite-arm cases (Shahriari et al.,, 2015). Moreover, Pryzant et al., (2023); Lin et al., (2023) use specific MAB methods targeting regret minimization to perform prompt selection, which, as further illustrated in Sec. 3.3, are not well-suited as they optimize the cumulative selection performance over a period instead of the final selection output. Thus, compared with this work, existing investigations either lack a comprehensive discussion of the connection between prompt optimization and MAB, or choose unsuitable MAB techniques to tackle prompt optimization. Moreover, as illustrated in Sec. 5, the TRIPLE solution outperform the previously adopted methods empirically.
2 Prompt Optimization under a Limited Budget
@@
Following Zhou et al., (2022); Chen et al., (2023), we present a concrete formulation of the problem of prompt optimization. Consider that we are using an LLM $f(\cdot)$, which provides a mapping from any input $X\in{\mathcal{V}}$ to a distribution $\Delta_{{\mathcal{V}}}$ over the language space ${\mathcal{V}}$. The answer $\hat{Y}\in{\mathcal{V}}$ given by the LLM is assumed to be sampled from $f(X)$ as $\hat{Y}\sim f(X)$. Note that instead of treating $f(\cdot)$ as a deterministic function providing a specific output answer, we generally consider the practical setting where the answers of LLM exhibit a certain level of randomness.
@@
For prompt optimization, we aim to find a prompt $p$ such that when concatenated with inputs $X$ of a certain task (i.e, as $[p;X]$), it provides good performance in expectation with respect to the input distribution ${\mathcal{I}}_{X}$ and the inherent randomness of LLM $f(\cdot)$. The performance is measured as $\mu(p):=\mathbb{E}_{X\sim{\mathcal{I}}_{X}}\mathbb{E}_{\hat{Y}\sim f([p;X])}[s(X,\hat{Y})]$, where $s(X,\hat{Y})$ denotes a score function that measures the quality of the output $\hat{Y}$ for the input $X$.
Motivated by the common usage scenario of LLMs, recent studies have imposed several constraints on this learning problem (Zhou et al.,, 2022; Pan et al.,, 2023; Chen et al.,, 2023; Guo et al.,, 2023), where the three key ones are (I) black-box: the method can be applied to black-box LLMs, i.e., only have access to an API $f(\cdot)$ and no access to the intermediate structure or parameters inside (including gradients, output likelihood, etc.); (II) discrete: the learned prompt must be discrete characters, instead of continuous values (i.e., soft prompts); and (III) interpretable: the learned prompt must be understandable by humans, instead of gibberish words.
@@
Intuitively, the process of learning a good prompt requires interactions with the LLM (i.e., sample $\hat{Y}\sim f([p;X])$ and evaluating its responses (i.e., obtain score $s(X,\hat{Y})$). However, as mentioned in Sec. 1, such interactions and evaluations are costly. Thus, besides the aforementioned constraints, we further explicitly take into account that the prompt optimization process should have (IV) a limited budget: the total number of trials with the LLM that happen during the learning is at most $N$. Finally, the prompt optimization problem considered in this work can be formulated as: finding $p^{*}$ with high performance $\mu(p^{*})$ under constraints of black-box, discrete, interpretable, and a limited budget.
@@
Directly tackling this prompt optimization problem has been widely recognized as challenging even without the constraint of a limited budget (Liu et al.,, 2023). As highlighted in Pryzant et al., (2023); Chen et al., (2023), it essentially requires performing a black-box discrete optimization. Instead, many proposed methods rely on the pipeline of first generating a pool of candidate prompts and then selecting from it (Jiang et al.,, 2020; Zhou et al.,, 2022; Xu et al.,, 2022; Prasad et al.,, 2022; Guo et al.,, 2023). The prompt generation can either be performed manually or follow designed automatic protocols. For example, the famous APE design (Zhou et al.,, 2022) selects from prompts generated by an LLM using demonstrations. From a unified perspective, we can simplify the problem into generating a pool of prompts ${\mathcal{P}}$ and finding the optimal prompt in it: $p^{*}:=\operatorname*{arg\,max}\nolimits_{p\in{\mathcal{P}}}\mu(p)$.
@@
Figure 1: The commonly adopted prompt optimization pipeline. Previous works mostly investigate the generation component and ignore costs during selection, where GrIPS and APE are proposed in Prasad et al., (2022); Zhou et al., (2022). This work, instead, focuses on the selection component under an explicit budget constraint.
While many efforts have been devoted along this line, we recognize that they are largely focused on how to generate prompts, while limited attention has been paid to how to select from the already generated prompts (as mentioned in Sec. 1 and further discussed in Appendix A). Naive treatments, such as uniformly evaluating all prompts, are understandable since budget limitations are not considered previously, i.e., unlimited evaluations can be performed. With an explicit budget limitation, however, we need to carefully allocate the budgets to each prompt so that the optimal prompt (or at least a sufficiently good one) can be correctly learned, which is the main focus of this work. An overview of the considered prompt optimization pipeline and our focus is illustrated in Fig. 1.
3 Connecting Prompt Optimization with Best Arm Identification
@@
We provide a new perspective of prompt optimization through the lens of tools in multi-armed bandits (MAB) (Lattimore and Szepesvári,, 2020; Lai and Robbins,, 1985). In particular, prompt optimization under a limited budget is shown to be intrinsically aligned with the problem of fixed-budget best-arm identification (BAI-FB) (Audibert et al.,, 2010; Karnin et al.,, 2013). In the following, a brief introduction to MAB is first provided. Then, the connection between prompt optimization (especially selection) and MAB (especially BAI-FB) is established. Based on this connection, we propose to fully leverage the rich toolbox from BAI-FB to perform efficient prompt optimization.
@@ 3.1 Multi-armed Bandits
The research of multi-armed bandits (MAB) has a long and rich history; see representative surveys of Lattimore and Szepesvári, (2020); Bubeck et al., (2012). The most basic form of MAB, i.e., the finite-armed stochastic bandits, considers a system with a set ${\mathcal{K}}$ finite arms (i.e., actions) that provide stochastic rewards when pulled. When interacting with the system, the agent can select one arm $k\in{\mathcal{K}}$ to pull at each time, and she receives a stochastic reward: $r_{k}\sim\texttt{dist}_{k}(\nu_{k})$, where $\texttt{dist}_{k}(\nu_{k})$ denotes the action $k$’s reward distribution with an unknown expectation $\nu_{k}$.
The learning objective of the agent in MAB can be roughly divided into two categories: (1) regret minimization, which maximizes the expected cumulative rewards collected by the agent (Auer et al.,, 2002; Audibert et al.,, 2009; Garivier and Cappé,, 2011); (2) best arm identification, which targets at outputting the best arm $k^{*}=\operatorname*{arg\,max}_{k\in{\mathcal{K}}}\nu_{k}$ (Audibert et al.,, 2010; Garivier and Kaufmann,, 2016; Jamieson and Nowak,, 2014). These two objectives often require different learning strategies. Regret minimization typically relies on a carefully designed balance between exploration (i.e., obtaining new information) and exploitation (i.e., collecting higher rewards based on the previous information). Best arm identification, on the other hand, is also called pure exploration as it only focuses on obtaining information to find the best arm. We here particularly note that although the designs targeting regret minimization often can converge to the optimal arm $k^{*}$ given a sufficient period of time, they are known to be inefficient for the objective of best arm identification in the MAB studies.
3.2 A Bandit View of Prompt Optimization
Based on the above introduction, it can be intuitively understood that the prompt optimization (especially, selection) problem can be mapped into an MAB setting: